OpenStudy (anonymous):

Determine whether the integral from 2 to infinity of (x^2+x+2)/(x^4+x^2-1) converges or diverges.

OpenStudy (anonymous):

Does \[\int\limits_{2}^{\infty}\frac{x^2+x+2}{x^4+x^2-1}dx\] converge or diverge?

OpenStudy (anonymous):

I think you have to integrate using partial fractions, then solve . If the answer is a number/0 then it converges, if you end up with infinity then it diverges. :)